OpenStudy (anonymous):

@Mathematics There are 2 things that make me happy, girls and money. P(G) is the probability that a girl goes out with me, and P(M) is the probability I receive money. P(G) = 0.3 P(M) = 0.1 P(H|G and M) = 1 P(H| not G and M) = 0.9 P(H| G and not M) = 0.9 P(H|not G and not M) = 0.1 find P(H), the probability that I am happy. @Mathematics

OpenStudy (anonymous):

0.399?

OpenStudy (anonymous):

how did you work out that answer?

OpenStudy (anonymous):

Is it right? I use P(A|B) = P(A^B)/P(B). And P(A) + P(~A) = 1. Also P(A) = P(A^B)+P(A^~B). Specifically, G^M v G^~M v ~G^M v ~G^~M = U (the whole universe).So H can be found by summing up its intersections with each of these events. Let me know if you need more explanation.